w is cube root of unity then
(1+w)^7=A+Bw
then A,B are

Grade:9

1 Answers

Arun Kumar IIT Delhi
10 years ago
Hello Student,

\\ (1+\omega)^7=(1+\omega)(1+\omega)^6 \\=>(1+\omega)^7=(1+\omega)(-\omega^2)^6 \\=1+\omega \\=>A,B=1,1

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty